# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v1

@@ 3.5.2. Impact of the Size of $\mathcal{S}$
In TraCS-D, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of the size of $\mathcal{S}$, meaning that the MSE of the perturbed direction, $\mathrm{MSE}[\varphi^{\prime}]$, remains constant regardless of the size of $\mathcal{S}$.  The distance space $\mathcal{D}_{r(\varphi)}$ depends on the specific direction $\varphi$ and the size of $\mathcal{S}$. In the worst case, $\mathcal{D}_{r(\varphi)}$ is the diagonal of $\mathcal{S}$. We have shown that the MSE of $\mathcal{M}_{-}$ is quadratic to the size of the distance space after linear mapping in Section 3.2.4. Therefore, when $|\mathcal{D}_{r(\varphi)}|$ increases linearly, the expected error of the perturbed distance, i.e. $\mathrm{E}[|r^{\prime}(\varphi)-r(\varphi)|]$, also increases linearly.  However, the error in the direction space $\mathcal{D}_{\varphi}$ affects the error in $\mathcal{D}_{r(\varphi)}$ in a non-linear way. Specifically, under the Euclidean distance metric, this effect is $|\mathcal{D}_{r(\varphi)}|\cdot 2\sin(|\varphi-\varphi^{\prime}|/2)$, which is the chord length of the arc between $\varphi$ and $\varphi^{\prime}$.

### v2

@@ 3.5.2 Impact of the Size of $\mathcal{S}$
In TraCS-D, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ is independent of the size of $\mathcal{S}$, meaning that the MSE of the perturbed direction, $\mathrm{MSE}[\varphi^{\prime}]$, remains constant regardless of the size of $\mathcal{S}$. The distance space $\mathcal{D}_{r(\varphi)}$ depends on the specific direction $\varphi$ and the size of $\mathcal{S}$. In the worst case, $\mathcal{D}_{r(\varphi)}$ is the diagonal of $\mathcal{S}$. We have shown that the MSE of $\mathcal{M}_{-}$ is quadratic to the size of the distance space after linear mapping in Section 3.2.4. Therefore, when $|\mathcal{D}_{r(\varphi)}|$ increases linearly, the expected error of the perturbed distance, i.e. $\mathrm{E}[|r^{\prime}(\varphi)-r(\varphi)|]$, also increases linearly. Nonetheless, the error in the direction space $\mathcal{D}_{\varphi}$ affects the error in $\mathcal{D}_{r(\varphi)}$ in a non-linear way. Specifically, under the Euclidean distance metric, this effect is $|\mathcal{D}_{r(\varphi)}|\cdot 2\sin(|\varphi-\varphi^{\prime}|/2)$, which is the chord length of the arc between $\varphi$ and $\varphi^{\prime}$.